OpenStudy (anonymous):

Find the GCF of 56 and 40.

OpenStudy (anonymous):

idk

OpenStudy (welshfella):

you can do this by finding the prime factors 40 = 2*2*2*5 56 = 2*2*2*7

OpenStudy (anonymous):

so 2?

OpenStudy (anonymous):

yep

OpenStudy (welshfella):

What numbers are common to both ?

OpenStudy (anonymous):

2

OpenStudy (welshfella):

not 2

OpenStudy (welshfella):

2*2*2 is common

OpenStudy (anonymous):

so 2

OpenStudy (welshfella):

the GCF is the GREATEST number which will divide into both numbers

OpenStudy (welshfella):

2*2*2 = ?

OpenStudy (anonymous):

is it 280?

OpenStudy (welshfella):